画像処理工学 2 単位 (選択)

Image Processing  カルンガル, ステファン · ギディンシ · 講師/ 知能情報工学科知能工学講座

【授業目的】視覚認識技術に不可欠な画像処理の基本的な手法を理解させる．【授業概要】画像処理工学は，医療画像処理，工業用画像処理，視覚パターン処

理等の基礎になるディジタル画像処理手法の基本的事項を修得するための講義であり，画像処理の基本概念， 2 値画像処理，画像の変換と強調，画像の復元，画像の特徴抽出，パターンマッチング，パターン分類，画像処理システム，および工業用画像処理への応用について講述する. なお，工業用画像処理については，その分野の専門家に集中講義を依頼することにより，広い最新情報を修得できるようにしている.

【キーワード】画像処理，パターン認識【先行科目】『線形システム解析』 (1.0)，『信号処理』 (1.0)，『マイクロプロセッ

サ』 (1.0)

【関連科目】『信号処理』 (0.5) 【履修要件】線形システム解析，信号処理工学及び演習，マイクロプロセッサを

履修しておくことが望ましい【履修上の注意】なし【到達目標】視覚情報処理技術に不可欠な画像処理技術の基本的な手法を理解し，

応用力をつける．【授業計画】

1. ディジタル画像処理の特徴，画像のデータ構造

2. ヒストグラム，画像処理アルゴリズムの形態，画像の表現，データ構造

3. 画像の 2 値化， 2 値画像の連結性と距離

4. 連結成分の変形操作，図形の形状特徴

5. 画像の変換と強調

6. 平滑化と雑音除去

7. 画像の復元，画像の再構成，幾何学的変換

8. エッジ検出，線の検出

9. 領域分割，テクスチャ解析

10. マルチスペクトル画像処理， 3 次元画像処理，動画像解析

11. パターンマッチング，教師付き分類，教師なし分類

12. 画像処理システム

13. 工業用画像処理 1(工業用画像処理の要点，位置，形状の認識)

14. 工業用画像処理 2(欠陥の認識，表面情報の認識)

15. 予備日 16. 定期試験【成績評価基準】出席 (10%)，れぽーと，演習，小テスト (50%), 期末試験 (40%). 【教科書】田村秀行編著「コンピュータ画像処理入門」【参考書】高木幹雄 · 下田陽久監修「画像解析ハンドブック」東京大学出版会【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215734 【対象学生】開講コースと同学科の夜間主コース学生も履修可能 (読) かいこう

こーすとどうがっかのがくせいもりしゅうかのう【連絡先】

⇒ カルンガル (088-656-7488, karunga@is.tokushima-u.ac.jp) ���� (オフィス

アワー: 8.30-5.30) 【備考】

⋄ 確率統計学，信号処理工学，線形システム解析および計算機アーキテクチャ

を履修しておくこと. ⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である

Image Processing 2 units (selection)

Stephen Karungaru·Githinji · Associate Professor / Intelligent Systems, Department of Information Science and Intelligent Systems

Target⟩ Understanding indispensable basic image processing methods for visual

recognition. Outline⟩ Image processing engineering is an entrance level lecture to master the

basics of digital image processing techniques in medical imaging, industrial images, vision pattern processing, etc. The lectures includes the fundamental concepts of binary image processing, image conversion, emphasis, restoration, feature abstraction, pattern matching, classification, application of image processing systems and industrial image processing. Computer laboratory exercises will be conducted quarterly. When possible, a lecture on the most recent treads in industrial image processing by a specialist in the field could also be included.

�������⟩ image processing, pattern recognition Fundamental Lecture⟩ “Linear System Analysis”(1.0), “Signal Processing”(1.0),

“Microprocessors”(1.0) Relational Lecture⟩ “Signal Processing”(0.5) Requirement⟩ Linear system analysis” (1.0), “Signal processing” (1.0), “Micro-

processors” (1.0) Notice⟩ None Goal⟩ indispensable basic image processing methods for visual recognition

Schedule⟩

1. Features of digital image processing and image data structures

2. Histogram

3. Image formation, binary images, distance

4. Transformation operations, connected elements, shape features

5. Image conversion and emphasis

6. Smoothing and noise elimination

7. Restoration, re-composition and geometrical conversion

8. Edge and line detection

9. Domain decomposition and texture analysis

10. 3D image processing and motion sequence analysis

11. Pattern matching and classification

12. Image processing system

13. Industrial image processing and shape recognition 14. Defects and surface information recognition

15. Questions and summary

16. Exam

Evaluation Criteria⟩ Result: attendance (10%) and reports, programming exercise

and continuous assessment tests (50%), final exam (40%). Textbook⟩ Hideyuki Tamura ”Introduction to computer image processing” Reference⟩ Mikio Takagi and Shimoda Yohisashi ”Image analysis handbook”

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215734

Student⟩ Can be taken by all students in the day and night courses

Contact⟩

⇒ Karungaru (+81-88-656-7488, karunga@is.tokushima-u.ac.jp) ���� (Office

Hour: 8.30-5.30) Note⟩

⋄ To pass this class and to fully understand each lecture, two hours each for preparation and review are necessary. ⋄ This course will be evaluated as indicated in the evaluation criteria above The final exam will cover the whole course